# FORTALESA: Fault-Tolerant Reconfigurable Systolic Array for DNN Inference

### v1

@@ V-A Fault propagation analysis for transient faults
$ts$ Cycle of a systolic array execution
@@
Fault in the input register of a systolic array. An output of the convolutional layer can be expressed as
@@
Then, an error added to the output of the convolutional layer equals
@@
$c_{out,f}^{end}$ =$=\left\{\begin{array}[]{rcl}t_{w}\cdot N&\mbox{if}&C_{out}\geqslant t_{w}\cdot N\\ C_{out}&\mbox{if}&C_{out}<t_{w}\cdot N\end{array}\right.$ (23)
@@
$\Box_{f}^{end}$ =$=\left\{\begin{array}[]{rcl}t_{a}\cdot N&\mbox{if}&W_{out}\cdot H_{out}\geqslant t_{a}\cdot N\\ W_{out}\cdot H_{out}&\mbox{if}&W_{out}\cdot H_{out}<t_{a}\cdot N\end{array}\right.$ (32)
@@
V-B Fault propagation analysis for permanent faults
@@
$c_{out,f(i)}^{end}$ =$=\left\{\begin{array}[]{rcl}i\cdot N&\mbox{if}&C_{out}\geqslant i\cdot N\\ C_{out}&\mbox{if}&C_{out}<i\cdot N\end{array}\right.\ \forall\ i\in\{1,...,T_{w}\}$ (39)
@@
V-C Assessment of the proposed architecture
Faults in the DMR mode with averaging. The effect of the fault correction on the error added to the output is defined by the affected PE, whether it is a main PE or a shadow. If the fault happens in the main PE, then fault correction works the following way:
$y_{1}^{corr}$ $=\frac{y_{1}+e+y_{1}}{2}=y_{1}+\frac{e}{2}=y_{1}+\frac{e}{2^{1}}$
$y_{2}^{corr}$ $=\frac{y_{2}+e/2+y_{2}}{2}=y_{2}+\frac{e}{4}=y_{2}+\frac{e}{2^{2}}$
$y_{n}^{corr}$ $=\frac{y_{n}+e/2^{n-1}+y_{2}}{2}=y_{n}+\frac{e}{2^{n}}$ (45)
where $y_{n}$ is the partial sum on the $n$-th cycle after the fault, $y^{corr}$ is the corrected partial sum used by the main PE. As the number of clock cycles after the fault increases, the value of the added error decreases approaching zero.
If the fault happens in the shadow PE, then fault correction works this way:
$y_{1}^{corr}$ $=\frac{y_{1}+y_{1}+e}{2}=y_{1}+\frac{e}{2}=y_{1}+\frac{e}{2^{1}}$
$y_{2}^{corr}$ $=\frac{y_{2}+e/2+y_{2}+e}{2}=y_{2}+\frac{3e}{4}=y_{2}+\frac{(2^{2}-1)e}{2^{2}}$
$\begin{split}y_{n}^{corr}&=\frac{y_{n}+(2^{n-1}-1)e/2^{n-1}+y_{2}+e}{2}\\ &=y_{n}+\frac{(2^{n}-1)e}{2^{n}}\end{split}$ (46)
As the number of clock cycles after the fault increases, the value of the added error approaches itself.
Faults in the DMR mode with setting mismatched bits to zero. Since the effect of setting mismatched bits to zero cannot be represented as an addition of an error term, values of affected outputs are re-calculated completely. Algorithm 1 presents an algorithm for calculating the effect of faults in the input register assuming the fault happens in the main PE.
Algorithm 1 An algorithm for calculating the effect of faults in the input register
  $c_{out,f}^{start}$ as in Eq. (20)
  $c_{out,f}^{end}$ as in Eq. (23)
  for $co=c_{out,f}^{start}$ to $c_{out,f}^{end}$ do
     $y_{0}\leftarrow 0$main PE}
     $y_{1}\leftarrow 0$shadow PE}
     for $ci=0$ to $C_{in}$ do
        for $i=0$ to $H_{k}$ do
           for $j=0$ to $W_{k}$ do
              if $ci=w_{cf}$ and $i=w_{if}$ and $j=w_{jf}$ then
                 $y_{0}\leftarrow y_{0}+x_{(ci,u+i,v+j)}\cdot w_{(co,ci,i,j)}+e$
                 $y_{1}\leftarrow y_{1}+x_{(ci,u+i,v+j)}\cdot w_{(co,ci,i,j)}$
              else
                 before the fault:
                 $y_{0}\leftarrow y_{0}+x_{(ci,u+i,v+j)}\cdot w_{(co,ci,i,j)}$
                 $y_{1}\leftarrow y_{1}+x_{(ci,u+i,v+j)}\cdot w_{(co,ci,i,j)}$
                 after the fault:
                 $y_{0}\leftarrow(y_{0}\&y_{1})+x_{(ci,u+i,v+j)}\cdot w_{(co,ci,i,j)}$
                 $y_{1}\leftarrow y_{1}+x_{(ci,u+i,v+j)}\cdot w_{(co,ci,i,j)}$
              end if
           end for
        end for
     end for
     $y_{(co,u,v)}=y_{0}$
  end for
@@ V-D Implementation
The workflow of the fault injection is presented in Fig. 7. Pytorch is used to run the inference of the network, leveraging GPU speed up. The fault injection module takes the output of the targeted layer and modifies only affected output values according to the aforementioned formulas. The simulation then continues with the erroneous values.
@@ VI Evaluation
VI-A Hardware parameters
The proposed architecture of the systolic array is described in SystemVerilog and synthesized using Cadence Genus 2021 and 45nm Nangate PDK. A comparison of the baseline systolic array without reconfiguration ability and four implementation options of FORTALESA is presented in Table IV. The size of the systolic arrays is 48$\times$48. Input and weight registers are 8-bit long, and output register is 32-bit long.
TABLE IV: Hardware implementation parameters
@@ VI-A Hardware parameters
PM-DMR0-TMR3 1.937 0.177 357
PM-DMR0-TMR4 1.929 0.176 372
PM-DMRA-TMR3 2.129 0.193 303
PM-DMRA-TMR4 2.091 0.190 302
VI-B Reliability analysis
Two common CNN models were used for reliability evaluation, including AlexNet, trained on the CIFAR-10 dataset, and VGG-11, trained on the ILSVRC-2012 dataset. Models and datasets were selected to represent different network scales. For the fault injection experiments, networks were quantized to 8-bit integer format. Experiments were performed on NVIDIA GeForce RTX 3090 24G GPU.
@@ VI-B Reliability analysis
For transient fault analysis, layer-wise statistical fault injection was done following the equation introduced in [42] to achieve 95% confidence and 5% error margin. For each fault injection, fault parameters (Table II) were set randomly and a test dataset of 10.000 inputs was fed to the network.
Fig. 8 and 9 show AVF values of each layer for selected networks considering different execution modes and configurations of the $48\times 48$ systolic array and transient faults. Faults were injected in each convolution layer using the proposed methodology. For TMR mode, it is assumed that all faults are corrected, and there are no output errors. Fully connected layers were not considered for fault injection as they occupy only one row of the systolic array and their vulnerability factor, therefore, is very low [23].
Results show the varying vulnerability of different layers and the masking effect of DMR mode implementations. It can be seen that the DMR mode decreases the vulnerability of certain layers almost twice. Those results can be used to decide how to assign different layers of DNNs to different execution modes of the systolic array.
Fig. 10 shows AVF values for reliability assessment of permanent faults for AlexNet. The analysis is done for stuck-at-1 faults since they are proved to be more critical [23]. Unlike transient faults, analysis is done for the whole network (all convolutional layers) as permanent faults persist throughout the network execution.
@@
VI-C Performance vs. reliability trade-off
A flexible performance vs. reliability trade-off can be achieved by executing different layers of the DNN using different modes of the reconfigurable systolic array. Fig. 11 and 12 plot reliability versus execution latency of the networks for all the possible combinations of modes. Latency is normalized to the execution of the network using performance mode for every layer (no fault tolerance). For reliability, Top1-class errors are considered. The Pareto front is marked with red color. Four implementations of FORTALESA are considered for each network since different implementations of DMR mode give slightly different reliability results and different implementations of TMR mode affect the latency.
(a) PM-DMR0-TMR3
(b) PM-DMR0-TMR4
(c) PM-DMRA-TMR3
(d) PM-DMRA-TMR4
Figure 11: Reliability vs. latency for AlexNet
(a) PM-DMR0-TMR3
(b) PM-DMR0-TMR4
(c) PM-DMRA-TMR3
(d) PM-DMRA-TMR4
Figure 12: Reliability vs. latency for VGG-11
To compare different implementations of FORTALESA, the AVF of the Pareto front points was plotted against a product of latency, power, area, and delay (Fig. 13 and 14). It can be seen that while systolic arrays with the TMR4 implementation have smaller area and power, and higher frequency parameters, taking execution latency into consideration version with the TMR3 implementation shows better overall results.
@@
VI-D Comparison with static redundancy
The comparison of the proposed run-time reconfigurable systolic array architecture with the static TMR approaches is presented in Fig. 15. A power-area product is plotted against the maximum possible throughput of the architecture (since the throughput of FORTALESA depends on the execution mode). Throughput is calculated as the number of MAC operations performed by the systolic array in one clock cycle multiplied by the frequency. Different cases are considered for static TMR: triplication of registers only, triplication of registers and MAC units, and triplication of the whole array. As well as different sizes: $48\times 48$ and $24\times 32$ as this is the efficient size of the $48\times 48$ systolic array operating in a TMR execution mode (TMR3 implementation option). As can be seen from the figure, the $24\times 32$ systolic array with static TMR has a lower power-area product than the proposed architecture but has a fixed low throughput. On the other hand, the $48\times 48$ systolic array with static TMR has high throughput, but the power-area product is significantly higher than the proposed architecture. FORTALESA, on average, requires 6×6$\times$6 × less resources than static TMR.
@@ VI-D Comparison with static redundancy
The proposed architecture is also compared with selective ECC proposed in [23]. While it protects only selected registers, it requires, on average, 2.5×2.5$\times$2.5 × more resources than FORTALESA, which protects all registers and MAC units.
VI-E Comparison with the state-of-the-art
The comparison of the proposed architecture with other methods to enhance fault tolerance of systolic arrays is presented in Table V. The proposed architecture protects both registers and MAC units in the PEs and since it utilizes a form of spatial redundancy it covers both transient and permanent faults. Faults are corrected in both fault-tolerant modes without interrupting inference execution for recovery.
TABLE V: Comparison with the state-of-the-art
@@ VII Conclusion
In this work, we proposed run-time reconfigurable fault-tolerant systolic array architecture FORTALESA with three execution modes and four implementation options. All four implementation options were evaluated in terms of resource utilization, throughput, and fault tolerance improvement. The proposed architecture is used for reliability enhancement of DNN inference on systolic array through heterogeneous mapping of different network layers to different execution modes. We also introduced a reliability assessment method based on fault propagation analysis for the evaluation of the proposed architecture and determination of the appropriate execution mode–layer mapping for DNN inference. The proposed architecture efficiently protects both registers and MAC units of systolic array PEs from transient and permanent faults without interrupting inference execution. The reconfigurability feature of FORTALESA enables a speedup up to 3×3$\times$3 ×, depending on layer vulnerability, and requires 6×6$\times$6 × less resources compared to static redundancy and 2.5×2.5$\times$2.5 × less resources than the previously proposed solution for transient faults.

### v2

@@ V-B Fault propagation analysis for transient faults
tsts Cycle of a systolic array execution
@@
An output of the convolutional layer can be expressed as
@@
A fault in a multiplier or register holding activation, weight, or partial sum results in an addition of an error $e$ to an output of the convolutional layer. This error $e$ depends on the error term $\varepsilon$ added to the faulty value and can be calculated for each type of fault from parameters given in Table II.
Fault in the input register of a systolic array. If the fault happens in the input register, it corrupts an activation, then an error added to the output of the convolutional layer equals
@@
$c_{out,f}^{end}$ $=\left\{\begin{array}[]{rcl}t_{w}\cdot N&\mbox{if}&C_{out}\geqslant t_{w}\cdot N\\ C_{out}&\mbox{if}&C_{out}<t_{w}\cdot N\end{array}\right$. (23)
@@
An error $e_{ireg}$ calculated using equation (14) is added to the affected output values.
@@
$\Box_{f}^{end}$ $=\left\{\begin{array}[]{rcl}t_{a}\cdot N&\mbox{if}&W_{out}\cdot H_{out}\geqslant t_{a}\cdot N\\ W_{out}\cdot H_{out}&\mbox{if}&W_{out}\cdot H_{out}<t_{a}\cdot N\end{array}\right$. (32)
@@
V-C Fault propagation analysis for permanent faults
@@
$c_{out,f(i)}^{end}$ $=\left\{\begin{array}[]{rcl}i\cdot N&\mbox{if}&C_{out}\geqslant i\cdot N\\ C_{out}&\mbox{if}&C_{out}<i\cdot N\end{array}\right.\ \forall\ i\in\{1,...,T_{w}\}$ (39)
@@ V-D Implementation
The workflow of the proposed reliability assessment method combining fault injection with fault propagation analysis is presented in Fig. 5. Pytorch is used to run the inference of the network, leveraging GPU speed up. The fault injection module takes the output of the targeted layer and modifies only affected output values according to the aforementioned formulas. The simulation then continues with the erroneous values. The implementation of the proposed reliability assessment method will be made open-source.
@@
Figure 8: Layer-wise reliability assessment of ResNet-18 considering transient faults
@@ VI Evaluation
VI-A Efficiency of reliability assessment method
The proposed reliability assessment method for systolic arrays is compared with the representative fault injection framework that models systolic array microarchitecture, Saca-FI [23]. While previous works rely on a cycle-accurate simulation of systolic array, which is a very time-consuming task, the proposed method is based on fault propagation analysis. Table IV presents the comparison of two fault injection methods. Runtime denotes the processing time of one input with fault injection. For runtime comparison, the selected model is VGG-16 evaluated on the ILSVRC-2012 dataset, systolic array size is $\times256$, fault is injected in the first convolutional layer. Both methods give AVF (Architecture Vulnerability Factor) as an output, the probability that a fault in the hardware structure causes an application output error [49]. While the functionality of the methods is similar, the proposed method is several orders of magnitude faster than the microarchitecture-level fault injection.
TABLE IV: Comparison of fault injection methods
Work Fault model Output Dataflow Runtime
Saca-FI Transient and permanent faults in registers AVF WS, IS, OS 47 s
Ours Transient and permanent faults in registers and multipliers AVF OS* 0.0028 s
* The proposed method can be easily extended to other dataflows as well.
VI-B FORTALESA hardware implementation parameters
The proposed architecture of the systolic array is described in SystemVerilog and synthesized using Cadence Genus 2021 and 45nm Nangate PDK. A comparison of the baseline systolic array without reconfiguration ability and four implementation options of FORTALESA is presented in Tables V and VI. Two sizes of the systolic array were selected: $\times48$ and $\times132$ to demonstrate the scalability of the design. Input and weight registers are 8-bit long, and the output register is 32-bit long. Due to the regular structure of the proposed architecture, the designs scale almost linearly proportionate to the increase in the number of processing elements. All implementations introduce around 5% overhead due to additional interconnects and MUXes, around 6-7% due to correction mechanisms for DRG0 versions, and around 17-18% for DRGA versions.
It can be noticed that frequency might drop considerably for the implementations with the averaging correction scheme for the DRG execution mode (DRGA). Critical path of these implementation variants includes an additional adder in the main PEs, thus increasing the datapath and lowering the frequency.
TABLE V: Hardware implementation parameters for $\times48$ systolic array
@@ VI-B FORTALESA hardware implementation parameters
PM-DRG0-TRG3 1.937 0.177 357
PM-DRG0-TRG4 1.929 0.176 372
PM-DRGA-TRG3 2.129 0.193 303
PM-DRGA-TRG4 2.091 0.190 302
TABLE VI: Hardware implementation parameters for $\times132$ systolic array
Implementations Area, mm2 Power, W Max frequency, MHz
Baseline SA 12.533 3.625 285
PM-DRG0-TRG3 14.243 4.053 252
PM-DRG0-TRG4 14.284 4.029 249
PM-DRGA-TRG3 15.826 4.552 230
PM-DRGA-TRG4 15.605 4.489 262
VI-C Reliability analysis
Experimental setup. Three common CNN models were used for reliability evaluation, including AlexNet, trained on the CIFAR-10 dataset, and VGG-11 and ResNet-18, trained on the ILSVRC-2012 dataset. Models and datasets were selected to represent different network scales. For the fault injection experiments, networks were quantized to an 8-bit integer format. Experiments were performed on NVIDIA GeForce RTX 3090 24G GPU.
@@ VI-C Reliability analysis
For transient fault analysis, layer-wise statistical fault injection was done following the equation introduced in [50] to achieve 95% confidence and 5% error margin. For each fault injection, fault parameters (Table II) were set randomly and a test dataset of 10.000 inputs was fed to the network. Faults were injected in each convolution layer using the proposed methodology (Section V). For the TRG execution mode, it is assumed that all faults are corrected, and there are no output errors. Fully connected layers were not considered for fault injection as they occupy only one row of the systolic array and their vulnerability factor, therefore, is very low [23].
Experimental results. Figures 6, 7, and 8 show AVF values of each layer for selected networks considering different execution modes and configurations of the $48\times 48$ systolic array and transient faults. For ResNet-18, only DRGA is considered since DRG0 correction scheme was not very efficient for this model.
@@
Results show the varying vulnerability of different layers and the masking effect of DRG execution mode implementations. It can be seen that the DRG execution mode decreases the vulnerability of certain layers almost twice. Those results can be used to decide how to assign different layers of DNNs to different execution modes of the systolic array.
Figure 9 shows AVF values for the reliability assessment of permanent faults for AlexNet. The analysis is done for stuck-at-1 faults since they are proven to be more critical [23]. Unlike transient faults, analysis is done for the whole network (all convolutional layers) as permanent faults persist throughout the network execution.
VI-D Performance vs. reliability trade-off
A flexible performance vs. reliability trade-off can be achieved by executing different layers of the DNN using different modes of the reconfigurable systolic array. Figures 10, 11 and 12 plot reliability versus execution latency of the networks and energy consumption for all the possible combinations of modes. Latency is normalized to the execution of the network using performance mode for every layer (no fault tolerance). Energy is given in milliwatt-hour (mWh). For reliability, Top1-class errors are considered. The Pareto front is marked with red color. Four implementations of FORTALESA are considered for each network since different implementations of DRG execution mode give slightly different reliability results, and different implementations of TRG execution mode affect the latency.
(a) PM-DRG0-TRG3
(b) PM-DRG0-TRG4
(c) PM-DRGA-TRG3
(d) PM-DRGA-TRG4
Figure 10: Reliability vs. latency and energy for AlexNet
(a) PM-DRG0-TRG3
(b) PM-DRG0-TRG4
(c) PM-DRGA-TRG3
(d) PM-DRGA-TRG4
Figure 11: Reliability vs. latency and energy for VGG-11
(a) PM-DRGA-TRG3
(b) PM-DRGA-TRG4
Figure 12: Reliability vs. latency and energy for ResNet-18
The left-most point of the Pareto front denotes the configuration where all layers are executed using performance mode. The execution latency is small, but there is no fault tolerance. The right-most point denotes the configuration where all layers are executed using TRG mode. The execution latency is high, but all faults are corrected. The points in between denote configurations with a combination of execution modes, including DRG mode. The figures show that DRG execution mode allows for a flexible reliability–performance trade-off.
To compare different implementations of FORTALESA, the AVF of the Pareto front points was plotted against a product of latency, power, area, and delay (Fig. 13 and 14). It can be seen that while systolic arrays with the TRG4 implementation have smaller area and power, and higher frequency parameters, taking execution latency into consideration version with the TRG3 implementation shows better overall results.
These figures present a design space exploration with the highlighted Pareto front, allowing designers to select a trade-off that fits the needs and requirements of their system.
@@
VI-E Comparison with static redundancy
The comparison of the proposed run-time reconfigurable systolic array architecture with the static TMR approaches is presented in Fig. 15. A power-area product is plotted against the maximum possible throughput of the architecture (since the throughput of FORTALESA depends on the execution mode). Throughput is calculated as the number of MAC operations performed by the systolic array in one clock cycle multiplied by the frequency. Different cases are considered for static TMR: triplication of registers only, triplication of registers and MAC units, and triplication of the whole array. As well as different sizes: $48\times 48$ and $24\times 32$ as this is the efficient size of the $48\times 48$ systolic array operating in a TRG execution mode (TRG3 implementation option). As can be seen from the figure, the $24\times 32$ systolic array with static TMR has a lower power-area product than the proposed architecture but has a fixed low throughput. On the other hand, the $48\times 48$ systolic array with static TMR has high throughput, but the power-area product is significantly higher than the proposed architecture. FORTALESA, on average, requires $6\times$ fewer resources than static TMR.
@@ VI-E Comparison with static redundancy
The proposed architecture is also compared with the selective ECC proposed in [23]. While it protects only selected registers, it requires, on average, $2.5\times$ more resources than FORTALESA, which protects all registers and MAC units.
VI-F Comparison with the state-of-the-art
The comparison of the proposed architecture with other methods to enhance fault tolerance of systolic arrays is presented in Table VII. The proposed architecture provides a comprehensive solution that protects both registers and MAC units in the PEs, and since it utilizes a form of spatial redundancy, it covers both transient and permanent faults. Faults are corrected in both fault-tolerant modes without interrupting inference execution for recovery. Additionally, the proposed architecture provides support for reconfigurability, addressing the dynamic requirements of AI applications.
TABLE VII: Comparison with the state-of-the-art
@@ VII Conclusion
In this work, we proposed a run-time reconfigurable fault-tolerant systolic array architecture FORTALESA with three execution modes and four implementation options. All four implementation options were evaluated in terms of resource utilization, throughput, and fault tolerance improvement. The proposed architecture is used for reliability enhancement of DNN inference on a systolic array through heterogeneous mapping of different network layers to different execution modes. We also introduced a reliability assessment method based on fault propagation analysis. It is used to efficiently evaluate the vulnerability of DNN models and select the appropriate execution mode–layer mapping for DNN inference. The proposed architecture efficiently protects both registers and MAC units of systolic array PEs from transient and permanent faults without interrupting inference execution. The reconfigurability feature of FORTALESA enables a speedup of up to $3\times$, depending on layer vulnerability, and requires $6\times$ fewer resources compared to static redundancy and $2.5\times$ fewer resources than the previously proposed solution for transient faults. As future work, an ASIC (Application-Specific Integrated Circuit) based on FORTALESA will be developed for further testing, e.g., at radiation facilities. The design of the chip will be made open-source.